Hobbes, the stuffed tiger, has a mass of 28.4 kg. Calvin, the little boy, has a mass of 17.7 kg. In a game of football, Hobbes runs to the right (towards Calvin) at 9.2 m/s. Calvin runs to the left (towards Hobbes) at 12.0 m/s. The two collide and stick together.



What is their final velocity? (Assume that velocities to the right are positive.)
1
Expert's answer
2020-04-15T10:37:23-0400

As per the given question,

mass of the Hobbes "(m)=28.4 kg"

mass of the little boy kelvin"(m_1)=17.7 kg"

Speed of the Hobbes (u)=9.2 m/sec

Speed of the kelvin "( u_1)=12 m\/sec"

They are running opposite to each other and after the collision, they are getting stick, let the final speed becomes V,

Now, applying the conservation of momentum,

"\Rightarrow 
mu+m_1u_1=(m+m_1)V"

Now, substituting the values, in the above,

"\Rightarrow V=\dfrac{28.4\times9.2-17.7\times 12}{28.4+17.7}" m/sec

"\Rightarrow V=\dfrac{48.88}{46.1}m/sec"

"V=1.06m\/sec"